Two motorcycles are travelling at speeds of 120 kilometers per hour and 90 kilometers per hour respectively. If the faster one travels for 4 hours longer than the slower one and also goes twice as far, how far did the slow one travel?
Answer by macston(5194)

.
T=time traveled by slower one.
.
(120km/hr)(T+4hrs)=2(90km/hr)(T)
120T+480=180T
480=(60T)
8=T
The slow one traveled 8 hours.
(8hrs)(90km/hr)=720km
ANSWER: The slower bike traveled 720 km
.
CHECK:
(120km/hr)(T+4hrs)=2(90km/hr)(T)
(120km/hr)(8hrs+4hrs)=2(90km/hr)(8hrs)
1440km=2(720km)
1440km=1440km